OpenStudy (anonymous):

The straight line l is parallel to y + 5x = 1, and passes through the point P(2, 3). Find its x-intercept

OpenStudy (baldymcgee6):

Well, we know that the slope of the parallel line is the same as the slope of the line of the given function

OpenStudy (anonymous):

5x?

OpenStudy (baldymcgee6):

we have y = -5x +1, right?

OpenStudy (baldymcgee6):

the general form of a line is y = mx+b, where m is the slope

OpenStudy (baldymcgee6):

do you get that?

OpenStudy (anonymous):

yes, so to find the x intercept do we have to plug it in to the form of a line?

OpenStudy (baldymcgee6):

first we need to find the equation of the parallel line, so now that we know the slope is m =-5, and they gave us the points (2,3)... y = mx+b 3 = -5(2) +b b = 13 y = mx+b y = -5x +13 x intercept is when y=0 0 = -5x +13 -13 = -5x x = 13/5 x int. is at (13/5, 0)

OpenStudy (anonymous):

oh alright, thanks for putting it in steps it really helps!

OpenStudy (baldymcgee6):

you're welcome